OpenStudy (anonymous):

Let f(x) = 3x2 + x − 3 and g(x) = x2 − 5x + 1. Find f(x) − g(x). 2x2 − 4x − 2 2x2 − 4x − 4 2x2 + 6x − 2 2x2 + 6x − 4

OpenStudy (anonymous):

@Mehek14 is it c bro ?

Mehek (mehek14):

\(3x^2-x^2=2x^2\)

Mehek (mehek14):

\(x-(-5x)=6x\)

Mehek (mehek14):

\(-3-1=-4\)

Mehek (mehek14):

so it should be D

OpenStudy (anonymous):

@Mehek14 is -3 extranous

Mehek (mehek14):

no